Suppose a business experiences a sudden increase in its fixed costs. For example, suppose property taxes...

  1. Suppose a business experiences a sudden increase in its fixed costs. For example, suppose property taxes increase dramatically. What impact, if any, will this have on the following:
    1. the firm's AFC (average fixed cost); will increase.
    2. the firm’s AVC (average variable cost); will be unaffected.
    3. the firm’s ATC (average total cost); will increase with the average variable cost.
    4. the firm’s MC (marginal cost); will increase.
  1. What changes, if any, is there likely to be in these same cost CURVES?
Answer #1

In the given case,

Total Fixed Cost i.e. TFC has increased. So,

a)

We know that

AFC=TFC/Q

Since there is an increase in TFC, AFC will increase at each output level. Statement is true.

b)

We know that

AVC=TVC/Q

Since there is no change in TVC. AVC will remain unaffected. Given statement is true.

c)

ATC=AFC+AVC

Since AFC will increase and AVC will remain the same, ATC will increase.

Given statement is false, as ATC will increase but AVC will remain the same. Rise is due to rise in AFC.

d)

Marginal Cost=Change in TVC/Change in Q

There is no change in TVC. So, MC will remain unaffected. Given statement is false

We can summarize as under

i) AFC will increase

ii) AVC will be unaffected.

iii) ATC will increase

iv) MC will be unaffected.